# On the Generalization of Handwritten Text Recognition Models

### v1

@@ IV-C Model selection and capacity
Model selection has no impact. Fig. 5 demonstrates that the choice of selection method has a negligible impact on OOD performance in the HTR models, with the results across the various selection strategies being practically identical. Contrary to findings in the broader (DG) literature [40], the oracle-based selection method does not yield substantial advantages over other strategies. These results, when applied to the oracle model, suggest that in the DG scenario, no “sweet spot” exists for selecting a model that consistently exhibits superior generalization performance.
More capacity is not useful. It can also be observed from Fig. 5 that an increase in model capacity (measured by the number of parameters) does not correspond to an improvement in the generalization capabilities in OOD scenarios, suggesting that increased model complexity does not necessarily translate to better generalization performance.
V Factor analysis of out-of-distribution
@@
Visual divergence. To quantify the visual differences between domains, we approached this problem as an Anomaly Detection (AD) task [91]. Following the reconstruction-based literature [92, 93, 94], we measure whether a sample of a target $T$ is out-of-distribution based on the reconstruction error of an Autoencoder (AE) trained on a source $S$ that we denote as $\phi_{\theta_{S}}$. In this work, we trained a simple convolutional AE ($\phi_{\theta_{S}}$) for each source dataset (details and results in Appendix IX), and we compute $\Delta(X_{S},X_{T})$ as the average reconstruction error (measured in MSE) as:
@@
Textual divergence. To quantify the divergence between two textual distributions ($Y_{S},Y_{T}$), we considered the averaged KL-divergence across varying n-grams [74, 95]. Specifically, we compute $\Delta(Y_{S},Y_{T})$ as the average KL-divergence for each n-gram from $n=1\dots 5$:
@@
where $V_{n}$ is the set of n-grams for the source vocabulary and $D_{KL}^{(n)}(P^{(S)}_{j}\parallel Q^{(T)}_{j})$ represents the KL divergence between the n-gram $j$ in the source and target distributions. Consequently, $\Delta(Y_{S},Y_{T})$ approximates how “unlikely” the text in the target domain is relative to the source domain. In our metrics, we denote $\Delta_{L}(Y_{S},Y_{T})$ as the divergence between the source text domain $Y_{S}$ and a synthetic target domain $T$ (the same ones used in the synthetic experiments) where the language $L$ is known. We refer to $\Delta_{GT}(Y_{S},Y_{T})$ as the actual divergence (see Appendix IX) between the source domain $Y_{S}$ and the ground truth of the target domain $Y_{T}$.
V-A Main latent factors
Thus, with all metrics collected (model parameters, visual divergence, textual divergence, ID and OOD errors and calibrations) and after applying standard preprocessing to normalize the columns, we extracted the $k$ eigenvectors corresponding to the $k$ eigenvalues $\geq 1$. In our case, $k=4$ (more details in Appendix X). After testing various rotations (although results did not differ significantly) the most interpretable explanation was achieved by oblimax rotation. Results are presented in Fig. 6, where color-coded (Pearson’s) correlations illustrate how each metric relates to the four hidden factors.
Factor 1: Textual divergence explains most of the OOD error. Factor 1 shows a strong positive correlation with textual metrics for both ground-truth $\Delta_{GT}(Y_{S},Y_{T})$ (0.92) and generic texts $\Delta_{L}(Y_{S},Y_{T})$ (0.84). This suggests that Factor 1 is heavily linked to textual divergences, highlighting it as a significant contributor to OOD error. Noteworthy, Factor 1 shows a moderate positive correlation with OOD error (0.62), implying that greater textual misalignment between domains leads to increased difficulty in generalizing for HTR models. Additionally, Factor 1 shows positive correlations OOD ECE (0.52), suggesting that it may also capture aspects of calibration linked to textual divergence, although to a lesser extent. Overall, these results underscore that alignment in textual content between source and target domain is crucial for reducing OOD error in HTR models.
@@ V-A Main latent factors
Factor 2: Visual similarity has a modest influence on OOD Error: Factor 2 has a high positive correlation with ID reconstruction error $\Delta_{S}(X_{S},X_{T})$ (0.8) and OOD reconstruction error $\Delta_{T}(X_{S},X_{T})$ (0.92) indicating that this factor predominantly reflects visual divergence between domains. Interestingly, Factor 2 also shows a weak to moderate negative correlation with OOD error (-0.3). This suggests that while visual divergence may not be a dominant factor in predicting OOD error, it does exert a modest influence. This relationship implies that models experience some performance degradation as the visual divergence between source and target domains increases, albeit this effect is weaker compared to the impact of textual divergence.
@@ VI Conclusions
VII Complementary results
In this section, we provide additional metrics to the Character Error Rate (CER) presented in the main paper. We provide the following information regarding the results: (1) Results from the main table of the paper (Table III) presented in terms of Word Error Rate (WER) in Table IV and (2) Dataset source for the best test result obtained for each architecture in Table V.
We observe in Table IV that the ID results remain generally consistent, comparable to those reported in the literature. However, the performance in the OOD scenario remains significantly poor. Additionally, the numerical values highlight that the WER, being a stricter metric than the CER, amplifies the performance gap in the OOD scenario. While the average ID to OOD gap was 37.6% in terms of CER, it increases to 60.3% when measured using WER. Table V presents the best-performing source domain for each target domain in the OOD scenario. The dominance of the IAM dataset is noticeable, accounting for nearly 60% of cases, followed by Bentham with slightly over 20%. However, as reported in Section IV-C, the choice of source domain has minimal impact on OOD performance, as all results remain poor even when the target domain is known (oracle scenario).
TABLE IV: In-distribution (ID) and out-of-distribution (OOD) results (WER %) for HTR models across datasets. The OOD result is reported from the best-performing source. Results marked with $\ast$ indicate outliers, meaning that the model did not converge in the ID setting. Average results (bottom row) are computed filtering out outliers. $\dagger$ denotes architectures implemented from the papers (no code provided).
@@ VII Complementary results
TABLE V: Best source domain for each target (rows) across all architectures studied in the paper.
@@ G.W. IAM Bentham IAM Bentham IAM IAM IAM IAM
VIII Hyperparameters
VIII-A Architectures implementation
As stated in the main paper, we aimed to follow the implementation closest to the original papers using the available information for those that did not provide code. In all cases, the most significant architectural change occurred in the final prediction layer, where the output vocabulary size was adjusted to match the vocabulary size (94) reported in Section III of the main text.
VIII-B Data augmentation
@@
Dilation details. The image is first inverted using cv2.bitwise_not. Then, cv2.dilate is applied with the selected kernel, expanding the white areas in the image. The process is repeated for the specified number of iterations. Finally, the image is inverted again to restore its original colors.
Erosion details. The image is inverted using cv2.bitwise_not, followed by cv2.erode with the selected kernel, shrinking the white areas. This operation is also repeated for the given number of iterations. Afterward, the image is inverted back to its original color scheme.
IX Visual and textual divergences
@@
IX-A Convolutional Autoencoder
IX-A1 Architecture
@@ IX-A2 Training details
IX-B Visual divergence
This section presents the results of visual domain divergence, measured by the reconstruction error obtained from the autoencoder described in previous sections. Fig. 8 illustrates the divergence (calculated as the average MSE per image) between each domain pair, with the source represented on the Y-axis and target on the X-axis. Divergences are computed between training and test splits for each pair. To facilitate interpretability, the values are normalized, such that a value of 100 reflects high divergence (lighter colors), while a value of 0 denotes indicating low divergence (darker colors). To validate the visual divergence results in the OOD scenario, Fig. 12 presents images from three pairs of domains with low visual divergence (left) and three domains with high one (right) with their respective scores. Note that the left column features writing styles with very similar stroke densities, while the right column displays styles that differ significantly in both stroke appearance and density. The domain pairs were selected based on the scores presented in Fig. 8, ensuring minimal repetition of domains to better highlight the differences.

### v2

@@ 4.3 Model selection and capacity
Model selection has no impact.
Fig. 5 demonstrates that the choice of selection method has a negligible impact on OOD performance in the HTR models, with the results across the various selection strategies being practically identical. Contrary to findings in the broader (DG) literature [33], the oracle-based selection method does not yield substantial advantages over other strategies. These results, when applied to the oracle model, suggest that in the DG scenario, no “sweet spot” exists for selecting a model that consistently exhibits superior generalization performance.
More capacity is not useful.
It can also be observed from Fig. 5 that an increase in model capacity (measured by the number of parameters) does not correspond to an improvement in the generalization capabilities in OOD scenarios, suggesting that increased model complexity does not necessarily translate to better generalization performance.
5 Factor analysis of out-of-distribution
@@
Visual divergence.
To quantify the visual differences between domains, we approached this problem as an Anomaly Detection (AD) task [93]. Following the reconstruction-based literature [15, 14, 53], we measure whether a sample of a target $T$ is out-of-distribution based on the reconstruction error of an Autoencoder (AE) trained on a source $S$ that we denote as $\phi_{\theta_{S}}$. In this work, we trained a simple convolutional AE ($\phi_{\theta_{S}}$) for each source dataset (details and results in Appendix 9), and we compute $\Delta(X_{S},X_{T})$ as the average reconstruction error (measured in MSE) as:
@@
Textual divergence.
To quantify the divergence between two textual distributions ($Y_{S},Y_{T}$), we considered the averaged KL-divergence across varying n-grams [62, 30]. Specifically, we compute $\Delta(Y_{S},Y_{T})$ as the average KL-divergence for each n-gram from $n=1\dots 5$:
@@
where $V_{n}$ is the set of n-grams for the source vocabulary and $D_{KL}^{(n)}(P^{(S)}_{j}\parallel Q^{(T)}_{j})$ represents the KL divergence between the n-gram $j$ in the source and target distributions. Consequently, $\Delta(Y_{S},Y_{T})$ approximates how “unlikely” the text in the target domain is relative to the source domain. In our metrics, we denote $\Delta_{L}(Y_{S},Y_{T})$ as the divergence between the source text domain $Y_{S}$ and a synthetic target domain $T$ (the same ones used in the synthetic experiments) where the language $L$ is known. We refer to $\Delta_{GT}(Y_{S},Y_{T})$ as the actual divergence (see Appendix 9) between the source domain $Y_{S}$ and the ground truth of the target domain $Y_{T}$.
5.1 Main latent factors
Thus, with all metrics collected (model parameters, visual divergence, textual divergence, ID and OOD errors and calibrations) and after applying standard preprocessing to normalize the columns, we extracted the $k$ eigenvectors corresponding to the $k$ eigenvalues $\geq 1$. In our case, $k=4$ (more details in Appendix 10). After testing various rotations (although results did not differ significantly) the most interpretable explanation was achieved by oblimax rotation. Results are presented in Fig. 6, where color-coded (Pearson’s) correlations illustrate how each metric relates to the four hidden factors.
Factor 1: Textual divergence explains most of the OOD error.
Factor 1 shows a strong positive correlation with textual metrics for both ground-truth $\Delta_{GT}(Y_{S},Y_{T})$ (0.92) and generic texts $\Delta_{L}(Y_{S},Y_{T})$ (0.84). This suggests that Factor 1 is heavily linked to textual divergences, highlighting it as a significant contributor to OOD error. Noteworthy, Factor 1 shows a moderate positive correlation with OOD error (0.62), implying that greater textual misalignment between domains leads to increased difficulty in generalizing for HTR models. Additionally, Factor 1 shows positive correlations OOD ECE (0.52), suggesting that it may also capture aspects of calibration linked to textual divergence, although to a lesser extent. Overall, these results underscore that alignment in textual content between source and target domain is crucial for reducing OOD error in HTR models.
@@ 5.1 Main latent factors
Factor 2: Visual similarity has a modest influence on OOD Error:
Factor 2 has a high positive correlation with ID reconstruction error $\Delta_{S}(X_{S},X_{T})$ (0.8) and OOD reconstruction error $\Delta_{T}(X_{S},X_{T})$ (0.92) indicating that this factor predominantly reflects visual divergence between domains. Interestingly, Factor 2 also shows a weak to moderate negative correlation with OOD error (-0.3). This suggests that while visual divergence may not be a dominant factor in predicting OOD error, it does exert a modest influence. This relationship implies that models experience some performance degradation as the visual divergence between source and target domains increases, albeit this effect is weaker compared to the impact of textual divergence.
@@ 6 Conclusions
Acknowledgments
This research was supported by the Spanish Ministry of Science and Innovation through the LEMUR research project (PID2023-148259NB-I00), funded by MCIU/AEI/10.13039/501100011033/FEDER, EU, and the European Social Fund Plus (FSE+).
The first author is supported by grant CIACIF/2021/465 from “Programa I+D+i de la Generalitat Valenciana”.
@@ Acknowledgments
7 Complementary results
In this section, we provide additional metrics to the Character Error Rate (CER) presented in the main paper. We provide the following information regarding the results: the main table of the paper (Table 3) is presented in terms of Word Error Rate (WER) in Table 4 and the dataset source for the best test result obtained for each architecture is reported in Table 5. Additionally, Tables 6 and 7 present all cross-domain results (both real and synthetic, respectively) for each architecture individually. In Table 6, the in-domain (ID) results are displayed on the main diagonals and highlighted in gray.
We observe in Table 4 that the ID results remain generally consistent, comparable to those reported in the literature. However, the performance in the OOD scenario remains significantly poor. Additionally, the numerical values highlight that the WER, being a stricter metric than the CER, amplifies the performance gap in the OOD scenario. While the average ID to OOD gap was 37.6% in terms of CER, it increases to 60.3% when measured using WER. Table 5 presents the best-performing source domain for each target domain in the OOD scenario. The dominance of the IAM dataset is noticeable, accounting for nearly 60% of cases, followed by Bentham with slightly over 20%. However, as reported in Section 4.3, the choice of source domain has minimal impact on OOD performance, as all results remain poor even when the target domain is known (oracle scenario). Additionally,
[…]
Table 4: In-distribution (ID) and out-of-distribution (OOD) results (WER %) for HTR models across datasets. The OOD result is reported from the best-performing source. Results marked with $\ast$ indicate outliers, meaning that the model did not converge in the ID setting. Average results (bottom row) are computed filtering out outliers. $\dagger$ denotes architectures implemented from the papers (no code provided).
@@ 7 Complementary results
Table 5: Best source domain for each target (rows) across all architectures studied in the paper.
@@ G.W. IAM Bentham IAM Bentham IAM IAM IAM IAM
Table 6: Complete CER results in all datasets using real data.
Method S/T IAM Rimes G.W. Bentham S.G. Rodrigo ICFHR2016
CRNN [64] IAM 6.4 25.0 31.1 25.3 45.5 40.9 86.2
Rimes 35.4 3.7 49.0 50.2 52.3 47.1 87.9
G.W. 55.6 61.5 8.2 59.2 69.3 66.2 100.0
Bentham 34.9 45.3 32.2 4.7 57.8 43.8 78.7
S.G. 77.7 74.5 89.3 78.0 7.2 52.8 100.0
Rodrigo 65.7 61.4 71.8 66.3 33.6 1.7 85.3
ICFHR2016 74.9 78.4 81.6 75.4 77.9 75.6 5.2
VAN [20] IAM 6.6 21.3 34.5 26.6 39.8 38.5 82.9
Rimes 28.6 5.6 46.1 45.0 47.2 43.7 88.4
G.W. 73.7 67.4 9.3 59.3 67.1 69.6 100.0
Bentham 37.2 41.7 32.0 7.4 49.4 38.6 75.3
S.G. 96.1 85.0 93.1 83.1 7.8 57.7 100.0
Rodrigo 76.5 70.7 78.2 68.1 41.1 2.3 87.3
ICFHR2016 70.8 74.8 76.4 67.8 72.1 71.4 7.5
C-SAN [26] IAM 28.6 29.8 49.8 38.9 50.2 46.6 90.9
Rimes 31.5 21.3 50.7 45.9 51.1 45.7 87.0
G.W. 60.7 60.2 32.0 63.2 68.4 66.8 96.9
Bentham 54.7 51.3 58.3 26.6 64.8 45.2 83.4
S.G. 75.3 72.5 81.8 75.5 39.8 50.8 90.2
Rodrigo 72.7 69.1 78.1 68.2 35.0 38.5 86.9
ICFHR2016 78.4 81.8 86.6 78.7 86.1 81.2 75.3
HTR-VT [47] IAM 5.8 28.3 40.0 33.3 44.2 38.5 86.1
Rimes 33.7 7.9 46.2 48.2 51.0 46.1 81.9
G.W. 70.1 73.3 34.9 76.3 79.2 76.9 85.9
Bentham 44.4 49.8 38.6 8.4 58.1 45.4 79.6
S.G. 78.7 78.2 89.8 78.1 17.1 60.1 100.0
Rodrigo 66.4 64.1 68.7 67.8 36.5 3.9 85.4
ICFHR2016 74.9 77.1 78.8 73.3 76.2 71.9 11.6
Kang [39] IAM 8.0 32.0 44.0 39.4 51.8 60.6 96.2
Rimes 42.1 5.7 65.5 63.5 62.1 65.2 92.6
G.W. 82.8 81.8 78.4 77.7 78.3 77.4 100.0
Bentham 53.4 59.7 46.6 8.5 80.8 71.1 95.5
S.G. 100.0 100.0 100.0 100.0 78.7 86.5 100.0
Rodrigo 81.7 78.6 85.9 78.6 61.1 2.6 95.9
ICFHR2016 74.7 76.5 75.3 74.1 75.9 74.9 7.8
Michael [54] IAM 7.5 35.5 43.6 43.5 55.3 65.3 85.2
Rimes 54.5 6.9 63.9 70.2 64.3 66.8 86.3
G.W. 78.9 80.5 53.8 73.4 79.5 76.7 100.0
Bentham 49.1 60.0 50.7 8.5 74.5 70.7 91.2
S.G. 100.0 100.0 100.0 100.0 76.9 79.5 100.0
Rodrigo 100.0 100.0 87.2 100.0 92.9 3.8 92.4
ICFHR2016 89.3 89.7 81.6 80.4 77.9 77.9 9.5
LT [10] IAM 7.9 30.8 32.3 33.8 51.4 48.4 98.1
Rimes 42.9 5.0 52.5 61.9 60.8 56.4 90.5
G.W. 83.5 85.8 79.6 87.9 75.0 73.3 100.0
Bentham 42.0 55.4 39.6 6.0 65.4 51.7 92.4
S.G. 86.6 82.1 95.6 84.9 12.5 65.9 96.7
Rodrigo 73.1 67.6 80.9 70.8 37.8 2.0 90.7
ICFHR2016 78.0 81.3 81.3 77.9 77.4 76.5 5.9
VLT [11] IAM 8.9 29.4 32.1 33.3 48.2 47.4 89.7
Rimes 44.9 5.1 56.4 63.2 62.4 59.1 96.4
G.W. 69.5 72.2 25.2 69.6 76.3 75.9 100.0
Bentham 41.3 53.7 43.7 6.1 65.8 53.5 85.1
S.G. 91.8 83.3 98.4 86.9 9.2 58.9 100.0
Rodrigo 71.7 67.2 80.2 72.0 38.7 2.2 90.7
ICFHR2016 78.3 80.3 80.8 81.7 79.5 81.8 6.0
Table 7: Complete CER results in all datasets using synthetic data.
Method S/T IAM Rimes G.W. Bentham S.G. Rodrigo ICFHR2016
CRNN [64] WIT-en 11.9 22.3 16.9 26.9 25.8 28.1 78.5
WIT-fr 19.5 17.0 26.3 31.7 26.3 27.2 78.1
WIT-es 20.0 22.7 27.4 32.9 27.2 22.2 79.4
WIT-la 20.7 24.6 27.4 34.5 21.4 27.4 79.0
WIT-de 20.4 25.3 27.3 32.1 28.8 28.3 77.6
VAN [20] WIT-en 16.9 24.3 25.8 26.1 26.3 28.4 75.0
WIT-fr 22.4 19.4 31.7 33.1 25.3 25.8 76.9
WIT-es 23.1 23.5 32.6 34.8 26.0 23.2 77.5
WIT-la 22.7 24.5 34.1 34.8 23.5 28.2 77.7
WIT-de 21.9 26.0 30.8 33.8 28.2 29.6 74.3
C-SAN [26] WIT-en 32.0 38.0 47.8 42.7 35.7 38.8 83.5
WIT-fr 35.7 35.9 47.4 46.5 36.4 40.4 81.5
WIT-es 36.2 38.0 46.4 47.4 35.0 37.3 82.7
WIT-la 36.6 38.8 49.6 48.0 35.8 40.7 83.4
WIT-de 34.8 38.6 48.9 46.0 36.6 40.5 81.1
HTR-VT [47] WIT-en 20.7 31.5 26.3 28.3 29.4 32.2 77.5
WIT-fr 27.5 26.6 34.2 38.1 29.3 32.1 78.1
WIT-es 28.2 30.3 35.2 38.6 30.1 26.6 77.3
WIT-la 28.9 33.1 35.4 39.6 27.8 30.6 78.2
WIT-de 29.2 34.1 36.6 39.2 33.0 34.2 76.7
Kang [39] WIT-en 28.7 44.5 45.1 51.3 40.0 46.1 85.3
WIT-fr 38.6 33.1 47.7 57.4 36.2 41.9 87.8
WIT-es 37.7 49.1 70.3 66.4 42.8 48.7 100.0
WIT-la 26.2 29.6 39.6 41.2 22.7 35.3 95.3
WIT-de 32.3 40.1 42.0 43.9 32.4 43.0 84.2
Michael [54] WIT-en 20.6 34.0 30.5 36.6 35.2 43.4 83.6
WIT-fr 32.9 25.2 42.8 52.4 36.8 41.1 83.5
WIT-es 33.8 33.9 45.0 54.0 36.6 33.4 85.2
WIT-la 37.7 39.5 47.3 58.5 30.7 45.6 82.9
WIT-de 39.1 44.1 48.4 61.7 40.3 49.7 79.8
LT [10] WIT-en 13.8 25.1 16.3 21.4 23.4 28.0 80.6
WIT-fr 22.5 18.9 26.2 35.5 24.1 27.7 79.8
WIT-es 23.3 25.0 27.1 38.0 22.5 24.8 81.4
WIT-la 24.0 26.4 27.9 39.3 22.5 28.5 83.0
WIT-de 23.1 27.6 24.9 36.3 26.0 28.2 78.8
VLT [11] WIT-en 15.3 26.7 19.0 25.0 23.5 29.1 80.6
WIT-fr 23.0 19.5 25.9 37.3 23.3 27.9 79.5
WIT-es 24.9 27.0 28.4 40.6 25.1 24.2 79.8
WIT-la 26.0 28.3 29.9 42.0 26.6 31.0 82.1
WIT-de 23.6 25.9 27.0 39.2 24.4 28.2 79.2
[…]
IAM 74.9 80.3 6.8 78.9 28.6
Rimes 93.5 81.5 27.2 80.4 21.3
G.W. 78.6 79.7 17.3 83.3 31.1
S.G. 80.4 82.5 44.1 87.5 25.3
Bentham 85.4 78.4 17.9 76.7 33.6
Rodrigo 76.4 81.2 38.1 86.2 38.5
ICFHR2016 95.3 87.2 92.6 88.7 75.3
8 Hyperparameters
8.1 Architectures implementation
As stated in the main paper, we aimed to follow the implementation closest to the original papers using the available information for those that did not provide code. In all cases, the most significant architectural change occurred in the final prediction layer, where the output vocabulary size was adjusted to match the vocabulary size (94) reported in Section 3 of the main text.
8.2 Data augmentation
@@
Dilation details.
The image is first inverted using cv2.bitwise_not. Then, cv2.dilate is applied with the selected kernel, expanding the white areas in the image. The process is repeated for the specified number of iterations. Finally, the image is inverted again to restore its original colors.
Erosion details.
The image is inverted using cv2.bitwise_not, followed by cv2.erode with the selected kernel, shrinking the white areas. This operation is also repeated for the given number of iterations. Afterward, the image is inverted back to its original color scheme.
9 Visual and textual divergences
@@
9.1 Convolutional Autoencoder
9.1.1 Architecture
@@ 9.1.2 Training details
9.2 Visual divergence
This section presents the results of visual domain divergence, measured by the reconstruction error obtained from the autoencoder described in previous sections. Fig. 8 illustrates the divergence (calculated as the average MSE per image) between each domain pair, with the source represented on the Y-axis and target on the X-axis. Divergences are computed between training and test splits for each pair. To facilitate interpretability, the values are normalized, such that a value of 100 reflects high divergence (darker colors), while a value of 0 denotes indicating low divergence (lighter colors). To validate the visual divergence results in the OOD scenario, Fig. 11 presents images from three pairs of domains with low visual divergence (left) and three domains with high one (right) with their respective scores. Note that the left column features writing styles with very similar stroke densities, while the right column displays styles that differ significantly in both stroke appearance and density. The domain pairs were selected based on the scores presented in Fig. 8, ensuring minimal repetition of domains to better highlight the differences.
@@ 9.2 Visual divergence
9.3 Textual divergence
We present the results of textual domain divergences, quantified as the averaged KL-divergence across n-grams as described in the main text. Fig.  9 shows the divergence between textual distributions across domains, with the source represented on the Y-axis and the target the X-axis. Divergences are computed between training and test splits for each pair. Fig. 10 presents the textual divergences between real source domains (Y-axis) and synthetic domains (X-axis) for each language. In this case, the divergence is calculated between the training split of the source domain and the training split of the synthetic data, as these are used to compute the n-grams and train the models in the synthetic experiments. Both figures display normalized values, where a value of 100 indicates maximum divergence (darker colors) and 0 minimum divergence between texts.
10 Factor analysis
The selection of the number of factors is a crucial criterion for analyzing the outcomes of the factor analysis. Note that the first $k$ factors span the subspace defined by the first $k$ eigenvectors of the data matrix. To determine the number of factors $(n)$, the simplest rule of thumb involves retaining all eigenvectors with eigenvalues $\geq 1$. This can be simply visualized by plotting the eigenvalues in descending order using a scree plot, as shown in Fig. 12. Based on this analysis, we decided to retain four factors as stated in the main text of the paper.